### v1

@@ Abstract
Tensor Robust Principal Component Analysis (TRPCA) holds a crucial position in machine learning and computer vision. It aims to recover underlying low-rank structures and characterizing the sparse structures of noise. Current approaches often encounter difficulties in accurately capturing the low-rank properties of tensors and balancing the trade-off between low-rank and sparse components, especially in a mixed-noise scenario. To address these challenges, we introduce a Bayesian framework for TRPCA, which integrates a low-rank tensor nuclear norm prior and a generalized sparsity-inducing prior. By embedding the proposed priors within the Bayesian framework, our method can automatically determine the optimal tensor nuclear norm and achieve a balance between the nuclear norm and sparse components. Furthermore, our method can be efficiently extended to the weighted tensor nuclear norm model. Experiments conducted on synthetic and real-world datasets demonstrate the effectiveness and superiority of our method compared to state-of-the-art approaches.
keywords:
Bayesian inference; tensor recovery; tensor nuclear norm; low rankness
{MSCcodes}
68Q25, 68R10, 68U05
@@ 1 Introduction
With data becoming ubiquitous from diverse fields and applications, data structures are becoming increasingly complex with higher dimensions. Tensor, a multidimensional array, is an efficient data structure with broad applications, including machine learning [38] and computer version [39]. Meanwhile, high-dimension data always lie near a low-dimensional manifold, which can be interpreted by their low-rankedness. In matrix processing, low-rank assumption allows two-dimensional data recovery from incomplete or corrupted data [10]. However, expanding the low-rank concept from matrices to tensors remains an unresolved challenge. A main challenge in tensor analysis is that the tensor rank is not well defined. Various definitions of tensor rank have been proposed. For example, the CANDECOMP/PARAFAC (CP) rank, as described in [27], is based on the CP decomposition [24] and identifies the smallest number of rank-one tensors needed to represent a tensor. The Tucker rank [13], which stems from the Tucker decomposition [43], consists of a vector where each component corresponds to the rank of a matrix obtained by unfolding the original tensor. Furthermore, developments in tensor singular value decomposition (t-SVD) [26] have led to the tensor multi-rank [13] and tubal rank [25], both of which are analogous to the matrix singular value decomposition (SVD).
@@
(1) $\min_{\mathcal{X}=\mathcal{L}+\mathcal{S}}\left\|\mathcal{L}\right\|_{*}+\lambda\left\|\mathcal{S}\right\|_{1}.$
where $\|\mathcal{L}\|_{*}$ is the tensor nuclear norm as the convex relaxation to the certain tensor rank. Note that minimizing the rank is an NP-hard problem. Various approximations have been proposed to approach different tensor ranks [22, 50, 36]. Here, $\|\mathcal{S}\|_{1}$ is the $\ell_{1}$ norm of sparsity, and $\lambda>0$ is the parameter used to balance low-rankedness and sparsity.
In the TRPCA model, we can further reformulate the equality constrain by a penalty term and turn the optimization model (1) into
(2) $\min_{\mathcal{S},\mathcal{L}}\tfrac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{S}-\mathcal{L}\|_{F}^{2}+\theta_{2}\|\mathcal{S}\|_{1}+\theta_{3}\|\mathcal{L}\|_{*},$
@@
(3) $\mathcal{X}=\mathcal{L}+\mathcal{S}+\mathcal{E},$
@@
The selection of the parameters in the model (1) and (2) is critical. Under the t-SVD framework, the optimal parameter for $\lambda$ is suggested in [32] for the tensor nuclear norm. Nevertheless, it cannot be extended to other forms of tensor low-rank regularization, such as the weighted tensor nuclear norm. This issue becomes more serious when dealing with models involving multiple parameters in (2). Traditional parameter selection methods, including the discrepancy principle [34], $L$-curve [18], GCV [17], and RWP [1, 6], are often customized to specific regularization formulations and need iterative minimizations, which makes it inadequate for our tensor recovery problem in (2).
@@
Despite these successes, the adoption of VBI in TRPCA remains limited. To our best knowledge, only [53] has explored VBI for TRPCA, employing a generalized sparsity-inducing prior. However, this method directly expresses the low-rank tensor as a t-product of two smaller factor tensors, presupposing the tubal rank, and models the sparse component $\mathcal{S}$ with independent Gaussian priors, which may not be optimal for sparse data. In contrast, we propose an approach that employs a tensor nuclear norm prior to $\mathcal{L}$, eliminating the need for predefined tensor ranks. For the sparse component $\mathcal{S}$, we adopt a Laplace prior, which better captures sparse structures. This reformulation enhances model flexibility, offering a more principled and less restrictive approach to tensor recovery, thereby mitigating limitations posed by prior assumptions on tensor ranks or sparsity patterns. The primary contributions of this work are succinctly summarized as:
@@ 2.1 Notations
The inner product between two tensors ${\mathcal{A}}$ and ${\mathcal{B}}$ in ${\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}}$ is defined as ${\langle\mathcal{A},\mathcal{B}\rangle=\sum_{k=1}^{n_{3}}\left\langle\mathcal{A}_{::k},\mathcal{B}_{::k}\right\rangle}$. The complex conjugate of ${\mathcal{A}}$, which takes the complex conjugate of each entry of ${\mathcal{A}}$, is denoted as conj$(\mathcal{A})$. The conjugate transpose of a tensor ${\mathcal{A}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is a tensor $\mathcal{A}^{\ast}$ obtained by conjugate transposing each of the frontal slices and then reversing the order of transposed frontal slices 2 through $n_{3}$. The tensor ${\ell_{1}}$-norm of $\mathcal{A}$ is defined as ${\|\mathcal{A}\|_{1}=\sum_{ijk}\left|a_{ijk}\right|}$, and the Frobenius norm as ${\|\mathcal{A}\|_{F}=}$ ${\sqrt{\sum_{ijl}\left|a_{ijk}\right|^{2}}}$.
@@ 2.2 T-product and t-SVD
(4) $\operatorname{unfold}(\mathcal{A})=\left[\begin{array}[]{c}\mathcal{A}_{::1}\\ \mathcal{A}_{::2}\\ \vdots\\ \mathcal{A}_{::n_{3}}\end{array}\right],\text{ fold }(\operatorname{unfold}(\mathcal{A}))=\mathcal{A},$
@@
Here unfold()($\cdot$)( ⋅ ) maps ${\mathcal{A}}$ to a matrix of size ${n_{1}n_{3}\times n_{2}}$ and fold()($\cdot$)( ⋅ ) is its inverse operator. We introduce the notation $\mathbf{A}:=\operatorname{bdiag}(\mathcal{A})$ to concisely represent the block diagonal matrix derived from the tensor $\mathcal{A}$. Here, $\operatorname{bdiag}(\cdot)$ designates the block diagonalization operator, with the $i$-th block corresponding to $\mathcal{A}_{::i}$.
Now, we focus on applying the Discrete Fourier Transformation (DFT) to tensors. We represent the tensor $\mathcal{A}$ transformed by DFT along its third (tubal) dimension as $\overline{\mathcal{A}}$.i This transformation is executed using the MATLAB command $\operatorname{fft}$, specifically performed as $\overline{\mathcal{A}}=\operatorname{fft}(\mathcal{A},[],3)$. Conversely, to revert the tensor to its original form from $\overline{\mathcal{A}}$, we use the inverse operation with $\mathcal{A}=\operatorname{ifft}(\overline{\mathcal{A}},[],3)$. We also introduce the notation $\overline{\mathbf{A}}:=\operatorname{bdiag}(\overline{\mathcal{A}})$ to represent the block diagonal matrix constructed from the tensor $\overline{\mathcal{A}}$. Next, we introduce the definition of t-product.
@@
(5) ${\mathcal{A}*\mathcal{B}=\operatorname{fold}(\operatorname{bcirc}(\mathcal{A})\cdot\operatorname{unfold}(\mathcal{B})),}$
@@
Using the t-product framework, we define the identity tensor $\mathcal{I}\in\mathbb{R}^{n\times n\times n_{3}}$ as a tensor with its first frontal slice being the $n\times n$ identity matrix, while all subsequent frontal slices consist entirely of zeros. It is clear that ${\mathcal{A}*\mathcal{I}=\mathcal{A}}$ and ${\mathcal{I}*\mathcal{A}=\mathcal{A}}$ given the appropriate dimensions. In addition, A tensor ${\mathcal{Q}\in}$ ${\mathbb{R}^{n\times n\times n_{3}}}$ is orthogonal if it satisfies $\mathcal{Q}^{*}*\mathcal{Q}=\mathcal{Q}*\mathcal{Q}^{*}=\mathcal{I}.$ Moreover, we call a tensor as $f$-diagonal if each of its frontal slices is a diagonal matrix. Next, we define the tensor singular value decomposition as below:
@@
(6) ${\mathcal{A}=\mathcal{U}*\mathcal{D}*\mathcal{V}^{*}},$
where ${\mathcal{U}\in\mathbb{R}^{n_{1}\times n_{1}\times n_{3}},\mathcal{V}\in\mathbb{R}^{n_{2}\times n_{2}\times n_{3}}}$ are orthogonal tensors, and ${\mathcal{D}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}}$ is an ${f}$-diagonal tensor.
@@
(tensor tubal rank [32]) For a tensor ${\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}}$, its tubal rank, denoted as $\operatorname{rank}_{t}(\mathcal{A})$, is defined as the number of nonzero singular tubes of $\mathcal{D}$, where $\mathcal{D}$ comes from the t-SVD of $\mathcal{A}$, i.e. $\mathcal{A}=\mathcal{U}*\mathcal{D}*\mathcal{V}^{*}$. In other words, one has
$\operatorname{rank}_{t}(\mathcal{A})=\#\{i,\mathcal{D}_{ii:}\neq 0\}.$
For tensor ${\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}}$ with tubal rank $r$, we also have skinny t-SVD similar as matrix. The skinny t-SVD of $\mathcal{A}$ is $\mathcal{A}=\mathcal{U}*\mathcal{D}*\mathcal{V}^{*}$, where $\mathcal{U}\in\mathbb{R}^{n_{1}\times r\times n_{3}},\mathcal{D}\in\mathbb{R}^{r\times r\times n_{3}},\mathcal{V}\in\mathbb{R}^{n_{2}\times r\times n_{3}}$, in which $\mathcal{U}^{*}*\mathcal{U}=\mathcal{I}$ and $\mathcal{V}^{*}*\mathcal{V}=\mathcal{I}$. Minimizing the tubal rank is an NP-hard problem; we introduce a tensor nuclear norm as a convex relaxation.
@@ 3.1 The likelihood
In (3), we assume the observed data $\mathcal{X}$ can be decomposed into three parts: $\mathcal{X},\mathcal{S},\mathcal{N}$. Note that $\mathcal{E}$’s elements as independent and identically distributed (i.i.d.) from a zero-mean normal distribution with precision $\theta_{1}$. Then we obtain the likelihood function $p(\mathcal{X}|\mathcal{S},\mathcal{L},\theta_{1})$ characterizes the probability of observing $\mathcal{X}$ conditioned on $\mathcal{L}$, $\mathcal{S}$, and $\theta_{1}$. By exploiting the properties of the normal distribution, the likelihood function is expressed as:
(7) $p(\mathcal{X}|\mathcal{S},\mathcal{L},\theta_{1})\propto\theta_{1}^{\frac{n}{2}}\exp\left(-\frac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{S}-\mathcal{L}\|_{F}^{2}\right),$
@@
where $C_{1}$ is a constant term that does not depend on $\mathcal{L}$, $\mathcal{S}$, or $\theta_{1}$ and can be ignored in inference procedure.
@@ 3.2 The prior distributions
In Bayesian inference, the selection of prior distributions is a fundamental step that shapes the posterior beliefs about the unknown parameters. These priors encode our prior knowledge or assumptions about the variables of interest. Here, we choose appropriate prior distributions for $\mathcal{L}$ and $\mathcal{S}$, which represent distinct latent variables with unique characteristics. We remark that the choice of these priors is informed by regularization terms.

### v2

@@ Abstract
Tensor Robust Principal Component Analysis (TRPCA) holds a crucial position in machine learning and computer vision. It aims to recover underlying low-rank structures and to characterize the sparse structures of noise. Current approaches often encounter difficulties in accurately capturing the low-rank properties of tensors and balancing the trade-off between low-rank and sparse components, especially in a mixed-noise scenario. To address these challenges, we introduce a Bayesian framework for TRPCA, which integrates a low-rank tensor nuclear norm prior and a generalized sparsity-inducing prior. By embedding the priors within the Bayesian framework, our method can automatically determine the optimal tensor nuclear norm and achieve a balance between the nuclear norm and sparse components. Furthermore, our method can be efficiently extended to the weighted tensor nuclear norm model. Experiments conducted on synthetic and real-world datasets demonstrate the effectiveness and superiority of our method compared to state-of-the-art approaches.
Key words: Bayesian inference; tensor recovery; tensor nuclear norm; low rankness
@@ 1 Introduction
With data becoming ubiquitous from diverse fields and applications, data structures are becoming increasingly complex with higher dimensions. Tensor, a multidimensional array, is an efficient data structure with broad applications, including machine learning [39] and computer vision [40]. Meanwhile, high-dimensional data always lie near a low-dimensional manifold, which can be interpreted by their low rank. In matrix processing, the low-rank assumption allows two-dimensional data recovery from incomplete or corrupted data [11]. However, expanding the low-rank concept from matrices to tensors remains an unresolved challenge. A main challenge in tensor analysis is that the tensor rank is not well defined. Various definitions of tensor rank have been proposed. For example, the CANDECOMP/PARAFAC (CP) rank, as described in [28], is based on the CP decomposition [25] and identifies the smallest number of rank-one tensors needed to represent a tensor. The Tucker rank [14], which stems from the Tucker decomposition [45], consists of a vector where each component corresponds to the rank of a matrix obtained by unfolding the original tensor. Furthermore, developments in tensor singular value decomposition (t-SVD) [27] have led to the tensor multi-rank [14] and tubal rank [26], both of which are analogous to the matrix singular value decomposition (SVD).
@@
$\min_{\mathcal{X}=\mathcal{L}+\mathcal{S}}\left\|\mathcal{L}\right\|_{*}+\lambda\left\|\mathcal{S}\right\|_{1}.$ (1)
where $\|\mathcal{L}\|_{*}$ is the tensor nuclear norm as the convex relaxation to a certain tensor rank. Note that minimizing the rank is an NP-hard problem. Various approximations have been proposed to approach different tensor ranks [23, 52, 37]. Here, $\|\mathcal{S}\|_{1}$ is the $\ell_{1}$ norm of sparsity, and $\lambda>0$ is the parameter used to balance low-rankedness and sparsity.
In the TRPCA model, we can further reformulate the equality constraint by a penalty term and turn the optimization model (1) into
$\min_{\mathcal{S},\mathcal{L}}\tfrac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{S}-\mathcal{L}\|_{F}^{2}+\theta_{2}\|\mathcal{S}\|_{1}+\theta_{3}\|\mathcal{L}\|_{*},$ (2)
@@
$\mathcal{X}=\mathcal{L}+\mathcal{S}+\mathcal{E},$ (3)
@@
The selection of the parameters in the model (1) and (2) is critical. Under the t-SVD framework, the optimal parameter for $\lambda$ in (1) is suggested in [33] for the tensor nuclear norm. Nevertheless, it cannot be extended to other forms of tensor low-rank regularization, such as the weighted tensor nuclear norm. This issue becomes more serious when dealing with models involving multiple parameters in (2). Traditional parameter selection methods, including the discrepancy principle [35], $L$-curve [19], GCV [18], and RWP [1, 6], are often customized to specific regularization formulations and need iterative minimizations, which makes it inadequate for our tensor recovery problem in (2).
@@
Despite these successes, the adoption of VBI in TRPCA remains limited. To our best knowledge, only [55] has explored VBI for TRPCA, employing a generalized sparsity-inducing prior. However, this method directly expresses the low-rank tensor as a t-product of two smaller factor tensors, presupposing the tubal rank, and models the sparse component $\mathcal{S}$ with independent Gaussian priors, which may not be optimal for sparse data. In contrast, we propose an approach that employs a tensor nuclear norm prior to $\mathcal{L}$, eliminating the need for predefined tensor ranks. For the sparse component $\mathcal{S}$, we adopt a Laplace prior, which better captures sparse structures. This reformulation enhances model flexibility, offering a more principled and less restrictive approach to tensor recovery, thereby mitigating limitations posed by prior assumptions on tensor ranks or sparsity patterns.
In comparison, joint maximum a posteriori (MAP) estimation minimizes the negative log posterior to obtain point estimates for $\mathcal{S}$, $\mathcal{L}$, and $\boldsymbol{\theta}$, simultaneously recovering tensors and parameters. Our VBI framework, however, approximates the full posterior distribution, enabling uncertainty quantification alongside point estimates. For practical applications such as denoising and background subtraction, we use the expectation of the variational distribution as the point estimate for $\mathcal{S}$ and $\mathcal{L}$, offering a robust and versatile approach to tensor recovery.
The primary contributions of this work are succinctly summarized as:
@@ 2.1 Notations
[…]
@@ 2.2 T-product and t-SVD
$\operatorname{unfold}(\mathcal{A})=\left[\begin{array}[]{c}\mathcal{A}_{::1}\\ \mathcal{A}_{::2}\\ \vdots\\ \mathcal{A}_{::n_{3}}\end{array}\right],\text{ fold }(\operatorname{unfold}(\mathcal{A}))=\mathcal{A},$ (4)
@@
Here unfold()($\cdot$) maps $\mathcal{A}$} to a matrix of size $n_{1}n_{3}\times n_{2}$} and fold()($\cdot$) is its inverse operator. We introduce the notation $\mathbf{A}:=\operatorname{bdiag}(\mathcal{A})$ to concisely represent the block diagonal matrix derived from the tensor $\mathcal{A}$. Here, $\operatorname{bdiag}(\cdot)$ designates the block diagonalization operator, with the $i$-th block corresponding to $\mathcal{A}_{::i}$.
Now, we focus on applying the Discrete Fourier Transformation (DFT) to tensors. We represent the tensor $\mathcal{A}$ transformed by DFT along its third (tubal) dimension as $\overline{\mathcal{A}}$. This transformation is executed using the MATLAB command $\operatorname{fft}$, specifically performed as $\overline{\mathcal{A}}=\operatorname{fft}(\mathcal{A},[],3)$. Conversely, to revert the tensor to its original form from $\overline{\mathcal{A}}$, we use the inverse operation with $\mathcal{A}=\operatorname{ifft}(\overline{\mathcal{A}},[],3)$. We also introduce the notation $\overline{\mathbf{A}}:=\operatorname{bdiag}(\overline{\mathcal{A}})$ to represent the block diagonal matrix constructed from the tensor $\overline{\mathcal{A}}$. Next, we introduce the definition of t-product.
@@
$\mathcal{A}*\mathcal{B}=\operatorname{fold}(\operatorname{bcirc}(\mathcal{A})\cdot\operatorname{unfold}(\mathcal{B})),$} (5)
@@
Using the t-product framework, we define the identity tensor $\mathcal{I}\in\mathbb{R}^{n\times n\times n_{3}}$ as a tensor with its first frontal slice being the $n\times n$ identity matrix, while all subsequent frontal slices consist entirely of zeros. It is clear that $\mathcal{A}*\mathcal{I}=\mathcal{A}$} and $\mathcal{I}*\mathcal{A}=\mathcal{A}$} given the appropriate dimensions. In addition, a tensor $\mathcal{H}\in$} $\mathbb{R}^{n\times n\times n_{3}}$} is orthogonal if it satisfies $\mathcal{H}^{*}*\mathcal{H}$$\mathcal{H}*\mathcal{H}^{*}=\mathcal{I}$. Moreover, we call a tensor $f$-diagonal if each of its frontal slices is a diagonal matrix. Next, we define the tensor singular value decomposition as below:
@@
$\mathcal{A}=\mathcal{U}*\mathcal{D}*\mathcal{V}^{*}$}, (6)
where $\mathcal{U}\in\mathbb{R}^{n_{1}\times n_{1}\times n_{3}},\mathcal{V}\in\mathbb{R}^{n_{2}\times n_{2}\times n_{3}}$} are orthogonal tensors, and $\mathcal{D}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$} is an f{f}-diagonal tensor.
@@
(tensor average rank and tubal rank [33]) The tensor average rank of $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$}, denoted as $\operatorname{rank}_{a}(\mathcal{A})$, is defined as
$\operatorname{rank}_{a}(\mathcal{A})=\frac{1}{n_{3}}\operatorname{rank}(\operatorname{bcirc}(\mathcal{A}))=\frac{1}{n_{3}}\sum_{i=1}^{n_{3}}\operatorname{rank}(\overline{\mathbf{A}}^{(i)}).$
The tensor tubal rank, denoted as $\operatorname{rank}_{t}(\mathcal{A})$, is defined as the number of nonzero singular tubes of $\mathcal{S}$, where $\mathcal{S}$ comes from the t-SVD of $\mathcal{A}$, i.e. $\mathcal{A}=\mathcal{U}*\mathcal{S}*\mathcal{V}^{*}$. In other words, one has
$\operatorname{rank}_{t}(\mathcal{A})=\#\{i,\mathcal{S}(i,i,:)\neq 0\}.$
For tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$} with tubal rank $r$, we also have skinny t-SVD similar as matrix. Minimizing the tubal rank is an NP-hard problem; we introduce a tensor nuclear norm as a convex relaxation.
@@
2.3 Probability distribution
Here, we define three kinds of probability distribution: the uniform distribution, the Gamma distribution, and the multivariate Gaussian distribution.
The uniform distribution is a distribution that assigns equal probability mass to a region. For $a,b\in R$ and $a<b$, the uniform distribution for a random variable $x\in R$ is defined as
$p(x)=\begin{cases}\frac{1}{b-a},&a\leq x\leq b,\\ 0,&\mathrm{otherwise}.\end{cases}$
The Gamma density function is given as
$p(x)=\mathcal{G}(x|a,b)\propto x^{a-1}\exp(-bx),$ (7)
where $a>0$ and $b>0$ represent shape and scale parameters respectively. We have its mean and variance of these Gamma distributions:
${\mathbb{E}(x)}=\frac{a}{b},\quad{\mathbb{V}\mathrm{ar}(x)}=\frac{a}{b^{2}}.$ (8)
The multivariate Gaussian distribution is fully characterized by a mean vector $\mu$ and a covariance matrix $\boldsymbol{\Sigma}$ and is defined as
$p({\boldsymbol{x}})=\mathcal{N}({\boldsymbol{x}}|\mu,\boldsymbol{\Sigma})=(2\pi)^{-\tfrac{n}{2}}|\boldsymbol{\Sigma}|^{-\tfrac{1}{2}}\exp\left(-\tfrac{1}{2}\left\|{\boldsymbol{x}}-\mu\right\|_{\boldsymbol{\Sigma}^{-1}}^{2}\right),$
where ${\boldsymbol{x}}\in\mathbb{R}^{n}$ is a random variable.
@@ 3.1 The likelihood
In (3), we assume the observed data $\mathcal{X}$ can be decomposed into three parts: $\mathcal{E},\mathcal{S},\mathcal{L}$. Note that $\mathcal{E}$’s elements as independent and identically distributed (i.i.d.) from a zero-mean normal distribution with precision $\theta_{1}$. Then we obtain the likelihood function $p(\mathcal{X}|\mathcal{S},\mathcal{L},\theta_{1})$ characterizes the probability of observing $\mathcal{X}$ conditioned on $\mathcal{S}$, $\mathcal{L}$, and $\theta_{1}$. By exploiting the properties of the normal distribution, the likelihood function is expressed as:
$p(\mathcal{X}|\mathcal{S},\mathcal{L},\theta_{1})\propto\theta_{1}^{\frac{n}{2}}\exp\left(-\frac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{S}-\mathcal{L}\|_{F}^{2}\right),$ (9)
@@
where $C_{1}$ is a constant term that does not depend on $\mathcal{S}$, $\mathcal{L}$, or $\theta_{1}$ and can be ignored in inference procedure.
@@ 3.2 The prior distributions
In Bayesian inference, the selection of prior distributions is a fundamental step that shapes the posterior beliefs about the unknown parameters. These priors encode our prior knowledge or assumptions about the variables of interest. Here, we choose appropriate prior distributions for $\mathcal{S}$ and $\mathcal{L}$, which represent distinct latent variables with unique characteristics. We remark that the choice of these priors is informed by regularization terms.